In the corner of a cellar is a pyramidal hoop of coal. The base of the hoop is an isosceles right triangle whose hypotenuse is 30ft. and its height is 8ft. What is the volume of the pyramidal hoop of coal in cu.ft.?


1
Expert's answer
2022-05-03T14:57:42-0400

isosceles right triangle:


"hypotenuse=30ft, leg_1=leg_2=\dfrac{30}{\sqrt{2}}ft"

"Area=A=\dfrac{1}{2}(\dfrac{30}{\sqrt{2}}ft)^2=225{ft}^2" The volume of the pyramidal hoop of coal is


"V=\dfrac{1}{3}(225{ft}^2)(8ft)=600{ft}^3"

600 cubic ft.